How many questions would I need to get right to get a 70% on a 76 question test
Answer by Alan3354(69443)

How many questions would I need to get right to get a 70% on a 76 question test
--------------
(70/100)*76 = 53.2 right
----
53/76 =~ 69.74%
54/76 =~ 71.05%
------------------
54, unless the % is rounded up.